(y-zx)p + (x+yz)q= x^2 + y^2


1
Expert's answer
2020-12-01T01:23:27-0500

P=y-zx, Q=x+yz, R=x2+y2

"\frac{dx}{y-zx}=\frac{dy}{x+yz}=\frac{dz}{x^2+y^2}"

Multiplyers: -y,-x,1

"\frac{-ydx-xdy+dz}{-y^2+xyz-x^2-xyz+x^2+y^2}=0"

-ydx-xdy+dz=0

-yx-xy+z=c1

c1=z-2xy

Multiplyers: y,x,-1

"\frac{ydx+xdy-dz}{y^2-zxy+x^2+zxy-x^2-y^2}=0"

ydx+xdy-dz=0

yx+xy-z=c2

c2=2xy-z

Answer: c1=z-2xy, c2=2xy-z